Find the length of the diagonal of a rectangle that has a length of 2ft and a width of 18in

Answer by tiffany222(56)

Convert 2 ft into inches:
2 ft = 24 in
Solve using the Pythagorean Theorem:
a^2 + b^2 = c^2
Let the length be a
Let the width be b
a^2 + b^2 = c^2
(24 in)^2 + (18 in)^2 = c^2
576 in^2 + 324 in^2 = c^2
900 in^2 = c^2
30 in = c
Final Answer: 30 in or 2.5 ft